a jet traveling 480 miles per hour leaves San Antonio for San Francisco, a distance of 1632 miles. An hour later another plane, going the same speed, leaves San Francisco for San Antonio. How long will it take before the planes to pass each other
Answer by richwmiller(17219)

In that 1 hour(s) the first plane has gone 480 miles
new distance 1152
(480+480)*t=1152
(960)*t=1152
t=1152/960
t=1.2 hours
1 hours
12.0 minutes
12 minutes rounded
0 seconds
time =1:12 hours to meet
time =1:12 pm rounded
The pass after 1:12 one hour 12 minutes
codedrttrt